Resolved -
The link is fully restored and services are stable. A post‑mortem will be published shortly.
Nov 28, 12:57 CET
Update -
The link has been fully restored, and the firewall cluster has recovered successfully. All systems are now operating normally, but we will keep monitoring them for the next few hours while we wait to receive a root‑cause explanation from the dark fibre vendor.
Nov 28, 11:23 CET
Update -
We are awaiting further information from our dark fibre vendor.
Nov 27, 16:40 CET
Update -
We identified the likely underlying cause. A flapping link on one of our dark fibre connections between data centre locations disrupted our firewall cluster as well as other network components. The link subsequently failed completely.
We are currently operating with reduced redundancy. An engineer from our dark fibre vendor is en route to verify the physical connection. We continue to monitor for any further issues.
Nov 27, 14:21 CET
Update -
We are continuing to monitor for any further issues.
Nov 27, 13:01 CET
Update -
The issue, caused by a network outage, has reoccurred. We are continuing to investigate this issue.
Nov 27, 12:00 CET
Update -
The issue, caused by a network outage, has reoccurred. We are continuing to investigate this issue.
Nov 27, 11:57 CET
Monitoring -
Connectivity has been restored, and all services are operational. We are currently investigating the root cause.
Nov 27, 11:29 CET
Update -
We are continuing to investigate this issue.
Nov 27, 11:21 CET
Investigating -
We are currently investigating this issue.
Nov 27, 11:19 CET